EPS and optimal debt ratio Williams Glassware has estimated, at various debt ratios, the expected earnings per share and theBusiness risk Financial risk Business risk Financial risk Luetuierl UI ValldLIUIT Coefficient of Variation 0.3+ 0.37
Answer #1

a]

The optimal debt ratio is the one where EPS is maximized

Maximum EPS appears to be at 60% debt ratio, with $3.93 per share earnings

b]

Coefficient of variation = standard deviation of EPS / EPS

Debt CV 1 ratio 2 0% 3 20% 4 40% 60% 6 80% EPS $ $ $ $ $ SD 2.33 $ 3.02 $ 3.52 $ 3.93 $ 3.82 $ 1.17 1.81 2.76 3.97 5.55 0.50

B C A Debt 1 ratio 2 0 3 0.2 4 0.4 5 0.6 6 0.8 EPS 2.33 3.02 3.52 3.93 3.82 SD 1.17 1.81 2.76 3.97 5.55 cv =C2/B2 =C3/B3 =C4/

c]

Graph 1 is correct.

As the debt ratio increases, the financial risk increases